# thanks: These two authors contributed equally to this workthanks: These two authors contributed equally to this work

### v1

@@ IV.2 Goldstone Mode
Figure 5: The log-log plot of $\chi_{k}$ versus system size $L$ for different inverse temperatures $\beta=1$ (red), 2 (blue), 4 (green), and 8 (black) at $\sigma=1.25$ (a), $1.75$ (b), and $1.875$ (c) is shown. As $L$ increases, the data points approach power-law growth with increasing system size $L$, confirming the form of correlation function Eq. (32), indicating the presence of a Goldstone mode. Furthermore, the fitting lines at different temperatures are parallel for a certain $\sigma$, aligning with the prediction that the value of $\eta_{\ell}$ remains independent of temperature.
As mentioned in the previous subsection, due to continuous symmetry breaking, we expect Goldstone mode in the LRO phase [47]. In this subsection, we demonstrate, both theoretically and numerically, the algebraic correlation function in the low-temperature phase for $\sigma<2$, i.e., $g(x)=g_{0}+c\cdot x^{-\eta_{\ell}}$ with $\eta_{\ell}=2-\sigma$ ($x$ and $g(x)$ represent the distance and correlation between two sites), induced by the Goldstone mode excitation. Moreover, we further speculate and investigate the possible logarithmic behavior of the system at $\sigma=2$.
Theoretical derivation. The reduced Hamiltonian of long-range O$(\mathcal{N})$ model can be written in momentum-space as [2]:
@@
where $\boldsymbol{q}$ denotes a $d$-dimensional momentum variable and $\boldsymbol{\Psi}(\boldsymbol{q})$ is the Fourier transform of a locally defined n-component spin field $\boldsymbol{\Psi}(\boldsymbol{x})$ ($\boldsymbol{\Psi}(\boldsymbol{x})$ and $\boldsymbol{\Psi}(\boldsymbol{q})$ respectively represent the spin field in real space and momentum space. For simplicity, they share the same symbol, and we hope this won’t lead to any ambiguities.); $t,K_{2},K_{\sigma}$, and $u$ are interacting parameters, and $t$ varies linearly with the distance to criticality.
In the last subsection, we have demonstrated the existence of spontaneous magnetization and continuous symmetry breaking at low temperatures for $\sigma\leq 2$. Thus, we can adapt mean-field approximation and consider small transverse fluctuations around it:
@@
Hence, for $\sigma<2$ at low temperatures, the correlation function takes on an algebraic form: $g(x)=g_{0}+c\cdot x^{-\eta_{\ell}}$, with $\eta_{\ell}=2-\sigma$. Note that the above derivation is applicable only at low temperatures; thus, we do not need to consider the renormalization of $K_{2}q^{2}$ terms as the system is far from criticality. Near or at the critical temperature, the system exhibits little to no spontaneous magnetization, and fluctuations play a significant role that cannot be considered negligible anymore. For $\sigma=2$, however, the derivation above is not applicable, and the exact form of the correlation function remains unknown. In Eq. (25), two terms—$\frac{K_{2}}{2}q^{2}$ and $K_{\sigma}q^{\sigma}$—become degenerate, potentially leading to logarithmic behaviors. Next, we present the numerical result to verify the form of the correlation function—$g(x)=g_{0}+c\cdot x^{-\eta_{\ell}}$—for $\sigma<2$, and the logarithmic behavior of the system at $\sigma=2$.
@@
With $\eta_{\ell}$ fixed at $2-\sigma$, the $\chi_{k}$ data can be fitted well, and the results are presented in Fig. 5. The figure displays $\chi_{k}$ as a function of system size $L$ in a double-logarithmic plot for $\sigma=1.25$, $1.75$, and $1.875$. In all cases, the data points for $\chi_{k}$ are close to a linear behavior, indicating the anticipated power-law relationship between $\chi_{k}$ and $L$. The slight upward curvature at smaller system sizes is attributed to finite-size correction terms, which become negligible for sufficiently large $L$. This linear trend in the large-$L$ regime demonstrates that the susceptibility follows the scaling relation $\chi_{k}\sim L^{2-\eta_{\ell}}$ and thus confirms the correlation function in Eq. (32). Additionally, the curves for different temperatures are almost parallel across all values of $\sigma$, and they can nearly be collapsed onto a single curve through vertical shifting. This observation indicates that the leading scaling exponent $\eta_{\ell}$ is temperature-independent. However, when we attempted to relax the value of $\eta_{\ell}$ and estimate it through fit using Eq. (33), the results exceeded our expected value of $2-\sigma$. This discrepancy is likely due to significant finite-size effects. To mitigate these large finite-size corrections, we further fit $(\xi/L)^{2}$ as a function of $L$ in the low temperature.
@@
In this approach, without considering any unknown correction terms, the data can be fitted well enough, indicating that $(\xi/L)^{2}$ indeed exhibits smaller finite-size correction. The detailed fitting results and final estimates for $\eta_{\ell}$ are presented in Table 1 and Table 3, respectively. The results in Table 3 are consistent with our conjecture that $\eta_{\ell}=2-\sigma$. While slight deviations are observed for $\sigma=1.75$ and $\sigma=1.875$ at $\beta=1$, these are reasonable given the proximity to criticality. Our theoretical derivation is most applicable in the low-temperature regime, where the system is far from the critical point, and fluctuations are minimal compared to spontaneous magnetization. At even lower temperatures, specifically for $\beta=2$, $4$, and $8$, all estimated values of $\eta_{\ell}$ are compatible with our conjecture.”
@@
$\chi_{k}=a_{0}L^{2}(\ln{L}+c_{1})^{\hat{\eta}_{\ell}},$ (35)
where $a_{0}$, $L_{0}$, and $\hat{\eta}_{\ell}$ are unknown parameters. The results of these fits are shown in Table 2. It is evident that, across different temperatures, the value of $\hat{\eta}_{\ell}$ from the fitting consistently approximates $-1$. So we fix $\hat{\eta}_{\ell}=-1$ and re-fit $\chi_{k}$ to Eq. (35), with results also presented in Table 2, supporting a possible scaling behavior: $\chi_{k}\sim L^{2}/\ln(L/L_{0})$.
In Fig. 6, we provide additional evidence by plotting $L^{2}/\chi_{k}$ as a function of $L$, with the $x$-axis in logarithmic coordinates. All the data points for different low temperatures show linear behavior, indicating $L^{2}/\chi_{k}\sim a\ln L+b$, which aligns with our assumption that $\chi_{k}\sim L^{2}/\ln(L/L_{0})$.
Figure 6: Demonstration of the logarithmic behavior for $\sigma=2$. $L^{2}/\chi_{k}$ is plotted in the semi-logarithmic coordinate as a function of system size $L$, with the $x$-axis in logarithmic scale, is shown for different inverse temperatures: $\beta=1$ (red), $\beta=2$ (blue), $\beta=4$ (green), and $\beta=8$ (black) at $\sigma=2$. The straight lines indicate the logarithmic behavior: $\chi_{k}\sim L^{2}/\ln(L/L_{0})$ for the $\sigma=2$ case.
@@ IV.3 $\sigma>2$ cases
It is well-established that for $\sigma>2$, the LR system belongs to the SR universality and exhibits quasi-long-range order (QLRO) at low temperatures. In this subsection, as a comparison to the $\sigma\leq 2$ case, we briefly demonstrate the low-$T$ properties of the QLRO phase for $\sigma>2$, using $\sigma=3$ and $\sigma=2.1$ as examples.
Figure 7 plots $\langle M^{2}\rangle$ with $L$ for $\sigma=3$ at double-log coordinates. The linearity of the data points indicates the scaling behavior $\langle M^{2}\rangle\sim L^{-\eta}$ and hence the correlation function’s form: $g(r)\sim r^{-\eta}$. As $L\rightarrow\infty$, $\langle M^{2}\rangle$ reduces to 0, thus no spontaneous magnetization at low temperature. Moreover, the slope of the fitted line for the data of $\beta=1$ is greater than that of $\beta=2$. This suggests that the value of $\eta$ decreases as the temperature approaches 0. All these properties are typical signatures of QLRO. The case of $\sigma=2.1$ is not presented in Fig. 7 because it exhibits strong finite-size corrections due to the adjacence to $\sigma=2$. As a result, $\langle M^{2}\rangle$ with $L$ forms a gradually flattening curve, making it difficult to determine whether the curve finally approaches a straight line as in QLRO or converges to a constant as in LRO.
Although it is difficult to study the power-law relationship between $\langle M^{2}\rangle$ and $L$ for $\sigma=2.1$, the QLRO nature of $\sigma=2.1$ can be revealed by analyzing the behavior of $(\xi/L)^{2}$. In Fig. 3, we have compared different behaviors of $(\xi/L)^{2}$ for varying $\sigma$. The diverging behavior for $\sigma\leq 2$ indicates the existence of LRO, whereas the converging behavior shows the nature of QLRO. For $\sigma=2.1$, due to significant finite-size corrections, $(\xi/L)^{2}$ converges slowly, causing some ambiguity. To clearly demonstrate the converging behavior of $(\xi/L)^{2}$, we perform a fit on its finite-size data. According to Eq. (10), at low temperatures with $L$ large enough, $(\xi/L)^{2}$ has the scaling: $(\xi/L)^{2}\sim\langle M^{2}\rangle/\langle M_{k}^{2}\rangle$. For $\sigma>2$, $\langle M^{2}\rangle$ and $\langle M_{k}^{2}\rangle$ share the same leading scaling exponent $\langle M^{2}\rangle\sim\langle M_{k}^{2}\rangle\sim L^{-\eta}$, but may have different correction terms. Therefore, $(\xi/L)^{2}$ should converge to a constant following a series of power-law decays. Thus $(\xi/L)^{2}$ can be fitted to:
@@
Figure 7: Demonstration of QLRO in the low-T phase for $\sigma=3$. The square of the magnetization density, $M^{2}$, is plotted against $L$ using double-logarithmic coordinates. For both $\beta=1$ and $\beta=2$, the linearity of the data points indicates a power-law relationship between $M^{2}$ and $L$, which is the signature of QLRO.
@@ IV.4 Revisiting the low-$T$ properties of the crossover point $\sigma=2$
Figure 9: The semi-logarithmic plot of $(\xi/L)^{2}/\ln(L/L_{0})$ as a function of $\sigma$ for several system sizes $L=64$ to 4096, with $y$-axis in logarithmic scale. The inset shows the data collapse of $(\xi/L)^{2}/\ln(L/L_{0}))$ vs. $\tilde{\sigma}\ln(L/L_{1})$, where $\tilde{\sigma}=\sigma-2$, $L_{0}=2.9$ and $L_{1}=3.0$.
The earlier subsections have identified the LRO for $\sigma\leq 2$ and the QLRO for $\sigma>2$ in the low-$T$ phase. To gain deeper insights into the low-$T$ properties of the system at the marginal point of $\sigma=2$ and the transition between LRO phase and QLRO phase, we fix the temperature at $T=1.0$, a value that is below the critical temperature $T_{c}(\sigma=2)=1.3671(4)$ but still sufficiently above the ground state, and observe the $2^{\text{nd}}$-moment correlation length $\xi$ as a function of $\sigma$ near $\sigma=2$. We first study the scaling form of $\xi$ at $\sigma=2$. Sec. IV.2 has clearly demonstrated the logarithmic scaling behavior of $\chi_{k}$: $\chi_{k}\sim L^{2}/\ln(L/L_{0})$ at $\sigma=2$ in the low-$T$ phase. Considering the defination Eq. (10) and the spontaneous magnetization in the low-$T$ phase, the $2^{\text{nd}}$-moment correlation length $\xi$ should exhibit similar logarithmic scaling form: $\xi\sim L\sqrt{\frac{\langle M^{2}\rangle}{\langle M_{k}^{2}\rangle}}\sim L\ln(L/L_{0})^{\frac{1}{2}}$. Next, by examining the behavior of $\xi$ as a function of $\sigma$ near $\sigma=2$, we can explore the transition between different scaling regimes and better understand how the system evolves as $\sigma$ varies.
Figure 9 presents a detailed finite-size scaling analysis of $(\xi/L)^{2}/\ln(L/L_{0})$ as a function of $\sigma$ for several system sizes $L$. Here, the constant $L_{0}=2.9$ according to the Table 2. A key feature of this figure is the clear crossing of curves at $\sigma=2$, marking a shift from LRO to QLRO. To capture the behavior near $\sigma=2$, we conjecture a refined scaling form for the $2^{\text{nd}}$-moment correlation length, taking into account logarithmic corrections at $\sigma=2$ in the low-temperature phase $T<T_{c}$. Specifically, we propose that $\xi$ scales as:
@@
where $\tilde{\sigma}=\sigma-2$ represents the deviation from $\sigma_{*}$, $\xi^{\prime}$ is a universal scaling function that governs the behavior near $\sigma=2$, and $L_{0}$ and $L_{1}$ are constants yet to be determined through data fitting. This scaling form accounts for the logarithmic corrections to the correlation length at the critical point, providing a more accurate description of the system’s behavior as $\sigma$ approaches $2$ from either side. The inset of Fig. 9 demonstrates the efficacy of this scaling hypothesis by showing a collapse of the scaled $\xi$ data points onto a single, universal curve. This data collapse provides numerical support for the conjectured scaling form and further confirms that $\sigma=2$ represents a marginal point where the system undergoes a transition between the LRO and QLRO phases at low temperatures.
V Results II: Power-law divergence of correlation length at high temperature

### v2

@@ IV.2 Goldstone Mode
Figure 5: The demonstration of the presence of Goldstone mode. The log-log plots of $\chi_{k}$ versus system size $L$ are shown for different temperatures at $\sigma=1.25$ (a), $1.75$ (b), and $1.875$ (c). As $L$ increases, the data points approach power-law growth with increasing system size $L$, confirming the form of the correlation function Eq. (31) and the presence of the Goldstone mode. Furthermore, the fitting lines at different temperatures are parallel for a certain $\sigma$, aligning with the prediction that the value of $\eta_{\ell}$ remains independent of temperature.
As discussed in the previous subsection, the spontaneous breaking of continuous symmetry leads to the emergence of Goldstone mode excitations in the LRO phase [51]. In this subsection, we provide both theoretical arguments and numerical evidence for the algebraic decay of the correlation function in the low-T phase for $\sigma<2$, i.e., $g(x)=g_{0}+c\cdot x^{-\eta_{\ell}}$ with $\eta_{\ell}=2-\sigma$. Here, $x$ and $g(x)$ represent the distance and correlation between two sites. This power-law behavior arises from the low-energy Goldstone mode excitations. We also examine the marginal case $\sigma=2$, where we conjecture a logarithmic decay of correlations and support this hypothesis through numerical analysis.
Theoretical derivation. The reduced Hamiltonian of long-range O$(n)$ spin model can be written in momentum-space as [2]:
@@
where $\boldsymbol{q}$ denotes a $d$-dimensional momentum variable and $\boldsymbol{\Psi}(\boldsymbol{q})$ is the Fourier transform of a locally defined $n$-component spin field $\boldsymbol{\Psi}(\boldsymbol{x})$ ($\boldsymbol{\Psi}(\boldsymbol{x})$ and $\boldsymbol{\Psi}(\boldsymbol{q})$ respectively represent the spin field in real space and momentum space. For simplicity, they share the same symbol, and we hope this won’t lead to any confusion); $t,K_{2},K_{\sigma}$, and $u$ are interacting parameters, and $t$ varies linearly with the distance to criticality.
In the last subsection, we have demonstrated the existence of spontaneous magnetization and continuous symmetry breaking at low temperatures for $\sigma\leq 2$. Thus, we can adapt the mean-field approximation and consider small transverse fluctuations around it:
@@
Hence, for $\sigma<2$ at low temperatures, the correlation function takes on an algebraic form: $g(x)=g_{0}+c\cdot x^{-\eta_{\ell}}$, with $\eta_{\ell}=2-\sigma$. Note that the above derivation is applicable only at low temperatures; thus, we do not need to consider the renormalization of $K_{2}q^{2}$ terms as the system is far from criticality. Near or at the critical temperature, the system exhibits little or no spontaneous magnetization, and fluctuations play a significant role that cannot be considered negligible anymore. For $\sigma=2$, however, the derivation above is not applicable, and the exact form of the correlation function remains unknown. In Eq. (24), two terms—$\frac{K_{2}}{2}q^{2}$ and $K_{\sigma}q^{\sigma}$—become degenerate, potentially leading to logarithmic behaviors. Next, we present the numerical result to verify the form of the correlation function—$g(x)=g_{0}+c\cdot x^{-\eta_{\ell}}$—for $\sigma<2$, and the logarithmic behavior of the system at $\sigma=2$.
@@
With $\eta_{\ell}$ fixed at $2-\sigma$, the $\chi_{k}$ data can be fitted well, and the results are presented in Fig. 5. The figure displays $\chi_{k}$ as a function of system size $L$ in a double-logarithmic plot for $\sigma=1.25$, $1.75$, and $1.875$. In all cases, the data points for $\chi_{k}$ are close to a linear behavior, indicating the anticipated power-law relationship between $\chi_{k}$ and $L$. The slight upward curvature at smaller system sizes is attributed to finite-size correction terms. This linear trend in the large-$L$ regime demonstrates that the susceptibility follows the scaling relation $\chi_{k}\sim L^{2-\eta_{\ell}}$ and thus confirms the correlation function in Eq. (31). Additionally, the curves for different temperatures are almost parallel across all values of $\sigma$, and they can nearly be collapsed onto a single curve through vertical shifting. This observation indicates that the leading scaling exponent $\eta_{\ell}$ is temperature-independent.
@@
In this approach, without considering any unknown correction terms, the data can be fitted well enough, indicating that $(\xi/L)^{2}$ indeed exhibits a smaller finite-size correction. The detailed fitting results and final estimates for $\eta_{\ell}$ are presented in Table 1 and Table 3, respectively. The results in Table 3 are consistent with our conjecture that $\eta_{\ell}=2-\sigma$. While slight deviations are observed for $\sigma=1.75$ and $\sigma=1.875$ at $\beta=1$, these are reasonable given the proximity to the critical point $\beta_{c}=0.68380(7)$ and $0.70737(7)$. Our theoretical derivation is most applicable in the low-T regime, where the system is far from the critical point, and fluctuations are minimal compared to spontaneous magnetization. At even lower temperatures, specifically for $\beta=2$, $4$, and $8$, all estimated values of $\eta_{\ell}$ are compatible with our derivation.
@@
$\chi_{k}=a_{0}L^{2}/(\ln{L}+c_{1})^{\hat{\eta}_{\ell}},$ (34)
where $a_{0}$, $L_{0}$, and $\hat{\eta}_{\ell}$ are unknown parameters. The results of these fits are shown in Table 2. It is evident that, across different temperatures, the value of $\hat{\eta}_{\ell}$ from the fitting consistently approximates $1$. So we fix $\hat{\eta}_{\ell}=1$ and re-fit $\chi_{k}$ to Eq. (34), with results also presented in Table 2, supporting a possible scaling behavior: $\chi_{k}\sim L^{2}/\ln(L/L_{0})$. This further implies that, at $\sigma=2$, the real-space correlation function asymptotically saturates to a non-zero constant as an inverse logarithm of distance, $g(x)\sim g_{0}+a\left(\frac{2\ln(x/b)-1}{\ln(x/b)^{2}}\right)$ for $x\gg 1$, where $a$ and $b$ are non-universal constants. This unusual inverse logarithmic decay of correlations highlights the nontrivial and marginal nature of the $\sigma=2$ case.
In Fig. 6, we provide additional evidence by plotting $L^{2}/\chi_{k}$ as a function of $L$, with the $x$-axis in logarithmic coordinates. All the data points for different low temperatures show linear behavior, indicating $L^{2}/\chi_{k}\sim a\ln L+b$, which aligns with our assumption that $\chi_{k}\sim L^{2}/\ln(L/L_{0})$. Moreover, from the fitting table 2, one can find that the amplitude $a_{0}$ is proportional to the temperature $T$. Therefore, we propose that $\chi_{k}$ should satisfy $\chi_{k}=A(T)L^{2}/\ln(L/L_{0})$ with $A(T)\propto T$ being an analytical function of $T$.
Figure 6: Demonstration of the logarithmic behavior for $\sigma=2$. $L^{2}/\chi_{k}$ is plotted in the semi-logarithmic coordinate as a function of the system size $L$ for different temperatures at $\sigma=2$. The linearity of data indicates the logarithmic behavior: $\chi_{k}\sim L^{2}/\ln(L/L_{0})$ in the low-T phase for the $\sigma=2$ case.
@@ IV.3 $\sigma>2$ cases
It is well-established that for $\sigma>2$, the LR system belongs to the SR universality and exhibits QLRO at low temperatures. In this subsection, as a comparison to the $\sigma\leq 2$ case, we briefly demonstrate the low-T properties of the QLRO phase for $\sigma>2$, using $\sigma=3$ and $\sigma=2.1$ as examples.
Figure 7 plots $\langle M^{2}\rangle$ as a function of $L$ for $\sigma=3$ at double-log coordinates. The linearity of the data points indicates the scaling behavior $\langle M^{2}\rangle\sim L^{-\eta}$ and hence the correlation function’s form: $g(r)\sim r^{-\eta}$. As $L\rightarrow\infty$, $\langle M^{2}\rangle$ reduces to 0, thus no spontaneous magnetization at low temperature. Moreover, the slope of the fitted line for the data of $\beta=1$ is greater than that of $\beta=2$. This suggests that the value of $\eta$ decreases as the temperature approaches 0. All these properties are typical signatures of QLRO. The case of $\sigma=2.1$ is not presented in Fig. 7 because it exhibits strong finite-size corrections due to the adjacency to $\sigma=2$.
The QLRO nature of $\sigma=2.1$ can be revealed by analyzing the behavior of $(\xi/L)^{2}$. In Fig. 3, we have compared different behaviors of $(\xi/L)^{2}$ for varying $\sigma$. The diverging behavior for $\sigma\leq 2$ indicates the existence of LRO, whereas the converging behavior shows the nature of QLRO. For $\sigma=2.1$, due to significant finite-size corrections, $(\xi/L)^{2}$ converges slowly, causing some ambiguity. To clearly demonstrate the converging behavior of $(\xi/L)^{2}$, we perform a fit on its finite-size data. According to Eq. (8), at low temperatures with $L$ large enough, $(\xi/L)^{2}$ has the scaling: $(\xi/L)^{2}\sim\langle M^{2}\rangle/\langle M_{k}^{2}\rangle$. For $\sigma>2$, $\langle M^{2}\rangle$ and $\langle M_{k}^{2}\rangle$ share the same leading scaling exponent $\langle M^{2}\rangle\sim\langle M_{k}^{2}\rangle\sim L^{-\eta}$, but may have different correction terms. Therefore, $(\xi/L)^{2}$ should converge to a constant following a series of power-law decays. Thus $(\xi/L)^{2}$ can be fitted to:
@@
Figure 7: Demonstration of QLRO in the low-T phase for $\sigma=3$. The squared magnetization, $\langle M^{2}\rangle$, is plotted against $L$ using double-logarithmic coordinates. For both $\beta=1$ and $\beta=2$, the linearity of the data points indicates a power-law relationship between $\langle M^{2}\rangle$ and $L$, which is the signature of QLRO.
@@ IV.4 Revisiting the low-T properties of the crossover point $\sigma=2$
Figure 9: The semi-logarithmic plot of $\chi_{k}L^{-2}\ln(L/L^{\prime}_{0})$ as a function of $\sigma$ for several system sizes $L=64$ to 4096, with $y$-axis in logarithmic scale. The inset shows the data collapse of $\chi_{k}L^{-2}\ln(L/L^{\prime}_{0})$ vs. $\tilde{\sigma}\ln(L/L_{1})$, where $\tilde{\sigma}=\sigma-2$, $L^{\prime}_{0}=2.89$ and $L_{1}=4.48$.
Figure 10: The semi-logarithmic plot of $(\xi/L)^{2}/\ln(L/L_{0})$ as a function of $\sigma$ for several system sizes $L=64$ to 4096, with $y$-axis in logarithmic scale. The inset shows the data collapse of $(\xi/L)^{2}/\ln(L/L_{0}))$ vs. $\tilde{\sigma}\ln(L/L_{1})$, where $\tilde{\sigma}=\sigma-2$, $L_{0}=0.75$ and $L_{1}=4.48$.
The preceding subsections have established the existence of LRO for $\sigma\leq 2$ and QLRO for $\sigma>2$ in the low-T phase. To gain deeper insights into the low-T properties of the system at the marginal point of $\sigma=2$, we investigate the $2^{\text{nd}}$-moment correlation length $\xi$ as a function of $\sigma$ near $\sigma=2$. We fix the temperature at $T=1.0$, a value that is below the critical temperature $T_{c}(\sigma=2)=1.3671(4)$. We begin by analyzing the scaling form of $\xi$ at $\sigma=2$. Section IV.2 has clearly demonstrated the logarithmic scaling behavior of $\chi_{k}$: $\chi_{k}\sim L^{2}/\ln(L/L_{0})$ at $\sigma=2$ in the low-T phase. Given the definition Eq. (8) and considering the finite magnetization in the low-T phase, the $2^{\text{nd}}$-moment correlation length $\xi$ should obey a similar logarithmic scaling form: $\xi\sim L\sqrt{\frac{\langle M^{2}\rangle}{\langle M_{k}^{2}\rangle}}\sim L\ln(L/L_{0})^{\frac{1}{2}}$. To explore the transition across $\sigma=2$, we analyze how the behavior of $\chi_{k}$ and $\xi$ as a function of $\sigma$.
Figure 9 and 10 presents a detailed finite-size scaling analysis of $\chi_{k}L^{-2}\ln(L/L^{\prime}_{0})$ and $(\xi/L)^{2}/\ln(L/L_{0})$ as a function of $\sigma$ for several system sizes $L$. A key feature is the clear crossing of curves at $\sigma=2$, signifying a change between LRO and QLRO. To capture the behavior near $\sigma=2$, we propose a refined scaling form for $\chi_{k}$ and $\xi$, incorporating logarithmic corrections at $\sigma=2$ in the low-T phase $T<T_{c}$. Specifically, we conjecture that $\chi_{k}$ scales as:
$\chi_{k}=L^{2}\ln(L/L^{\prime}_{0})^{-1}\chi_{k}^{\prime}(\tilde{\sigma}\ln\left(L/L_{1}\right)),$ (36)
and $\xi$ scales as:
@@
where $\tilde{\sigma}=\sigma-2$ represents the deviation from $\sigma_{*}$; $\chi^{\prime}_{k}$ and $\xi^{\prime}$ are universal scaling functions that govern the behavior near $\sigma=2$; $L_{0}$, $L^{\prime}_{0}$ and $L_{1}$ are constants determined via fitting. The insets of Fig. 9 and 10 demonstrate the efficacy of the scaling hypotheses by respectively showing a collapse of the scaled $\chi_{k}$ and $\xi$ data points onto single, universal curves. The data collapse provides strong numerical support for conjectured scaling forms and further confirms that $\sigma=2$ represents a marginal point where the system undergoes a transition between the LRO and QLRO phases at low temperatures.
V Power-law divergence of correlation length at high temperature